# Lemmanaid: Neuro-Symbolic Lemma Conjecturing

### v3

@@ Abstract
Automatically conjecturing useful, interesting and novel lemmas would greatly improve automated reasoning tools and lower the bar for formalizing mathematics in proof assistants. It is however a very challenging task for both neural and symbolic approaches. We present the first steps towards a practical neuro-symbolic lemma conjecturing tool, Lemmanaid, that combines Large Language Models (LLMs) and symbolic methods, and evaluate it on proof libraries for the Isabelle proof assistant. We train an LLM to generate lemma templates that describe the shape of a lemma, and use symbolic methods to fill in the details. We compare Lemmanaid against an LLM trained to generate complete lemma statements as well as previous fully symbolic conjecturing methods. Lemmanaid outperforms both neural and symbolic methods on test sets from Isabelle’s HOL library and from its Archive of Formal Proofs, discovering between 29-39.5% of the gold standard human written lemmas. This is 8-15% more lemmas than the neural-only method. By leveraging the best of both symbolic and neural methods we can generate useful lemmas for a wide range of input domains, facilitating computer-assisted theory development and formalization.
conjecturing, theorem proving, formalization, neuro-symbolic, proof assistants, interactive theorem proving, AI for math, lemma synthesis, theory exploration
[…]
@@ 1 Introduction
Learning to construct new, interesting, and useful lemmas for proof assistants is an important yet underexplored area in AI for mathematical reasoning (Yang et al., 2024). Such lemmas can aid a human user working on a mathematical formalization, as well as strengthen automated theorem provers. In this work, we examine how LLMs can be used for lemma generation, and how they can be combined with symbolic tools for optimal results. Our aim is to to provide a first tool towards generic conjecturing over a broad range of mathematical theories, which is practically useful for users of proof assistants.
A weakness of LLMs is that they sometimes generate repetitive or redundant lemmas, fail to discover more novel and useful lemmas, or hallucinate undefined symbols in the formalization. Furthermore, there are no correctness guarantees on the LLM’s output, so the generated lemmas may simply be false. These challenges have been encountered in previous work on neural conjecturing (Urban & Jakubův, 2020; Rabe et al., 2021; Johansson & Smallbone, 2023). Symbolic methods, on the other hand, can be designed and programmed to avoid repetition and redundancy. However, symbolic methods will only generate lemmas that fit a predefined specific search space, and tend to scale poorly to a larger search space. Previous symbolic tools  (Smallbone et al., 2017; Einarsdóttir et al., 2021; Singher & Itzhaky, 2021) have been used to successfully discover, for example, lemmas needed in automated (co-)inductive provers (Johansson et al., 2014; Einarsdóttir et al., 2018, 2024; Kurashige et al., 2024). However, these tools are limited in the shape, size and domain of lemmas they can generate, and do not scale well to larger sets of inputs.
@@
To address these shortcomings, we propose a novel neuro-symbolic lemma conjecturing approach and tool: Lemmanaid. An LLM is trained to generate lemma templates that describe the shape of a family of analogous lemmas, rather than directly generating complete lemmas. Symbolic synthesis methods are then used to fill in the details. Previous work has shown that such families of analogous lemmas indeed occur in proof assistant libraries (Einarsdóttir et al., 2022; Heras et al., 2013). In this way, we leverage the best of both neural and symbolic methods. The LLM suggests appropriate analogous lemma-patterns likely to be relevant for the theory at hand. The symbolic engine ensures correctness and novelty, while keeping the search space manageable. As far as we are aware, this is the first work focusing on neuro-symbolic lemma conjecturing.
For evaluation, we measure the coverage of (unseen test-set) lemmas that can be discovered by Lemmanaid from Isabelle’s HOL library [^1] https://isabelle.in.tum.de/dist/library/HOL/index.html and from its Archive of Formal Proofs (AFP) [^2] https://www.isa-afp.org. Using human-written formalizations as a gold-standard gives us a good approximation of how many interesting lemmas Lemmanaid can produce for a novel theory (taking only basic definitions of the theory as input), an evaluation strategy used in many prior lemma conjecturing works, e.g. (Montano-Rivas et al., 2012; Smallbone et al., 2017; Einarsdóttir et al., 2018; Urban & Jakubův, 2020). In our experiments, we use a DeepSeek-coder-1.3b language model. We choose a small LLM for several reasons: in addition to computational and environmental concerns, our ultimate aim is to create something that is accessible for regular proof-assistant users without need for huge compute resources, i.e. a model which can run locally for inference. We further show that the Lemmanaid approach compares favorably to the results achievable using purely neural or purely symbolic conjecturing.
The main contributions of our work are:
@@
  • Lemmanaid, the first neuro-symbolic lemma conjecturing approach that uses an LLM to predict templates and a symbolic engine to instantiate templates as lemmas.
@@
  • An evaluation of Lemmanaid on the Isabelle proof assistant’s HOL and AFP libraries. This goes much beyond evaluations of prior tools that have focused on specific domains.
@@
To ensure reproducibility of our results and enable others to build on our work, we make all code, experimental scripts, data, and models publicly available online [^3] https://anonymous.4open.science/r/Lemmanaid.
@@ 2 Related Work
Proof Assistants, such as Isabelle (Nipkow et al., 2002) and Lean (de Moura et al., 2015) are increasingly being used to check proofs in both mathematics and computer science for correctness. To do so, the user needs to formalize their theory, by translating it into the formal language of the proof assistant. They then interact with the system to construct a proof by stringing together calls to tactics, each executing and checking some part of the proof. Popular tools like Sledgehammer can automate (parts of) many proofs by selecting a suitable set of previously proved lemmas, and sending the conjecture to an automated first-order prover or SMT-solver (Blanchette et al., 2011).
Autoformalization and Proof Synthesis
@@
Formalizing theories in proof assistants is a non-trivial task, which has sparked interest in autoformalization: translating definitions, theorems and lemmas written in natural language to the formal language of a proof assistant (Wang et al., 2018; Szegedy, 2020; Wu et al., 2022). Furthermore, even with definitions and statements formalized, constructing the required proofs from various tactics is again non-trivial, even with the help of tools like Sledgehammer. This has motivated work on various LLM-driven methods for synthesizing proof scripts for proof assistants such as Isabelle or Lean, either incrementally (Jiang et al., 2023; Ren et al., 2025), or by generating whole proofs at once (First et al., 2023). The LEGO-prover furthermore attempts to introduce intermediate proof statements as reusable lemmas (Wang et al., 2024). Our work does not focus on producing proofs, but on suggesting suitable conjectures similar to the kind of lemmas that humans write down in libraries of formal proofs. As such it is orthogonal to much previous work, but could in the future be combined with, and complement, proof synthesis systems.
@@
Following the observation that many mathematical theories share analogical lemmas of similar shapes, Buchberger et al. (2006) proposed to use templates as guidance in mathematical theory exploration, allowing efficient conjecturing of many (but not all) lemmas by analogy to known shapes. This has been implemented in a range of symbolic lemma conjecturing systems, including some targeting Isabelle/HOL (Montano-Rivas et al., 2012; Heras et al., 2013; McCasland et al., 2017; Einarsdóttir et al., 2021; Nagashima et al., 2023). A similar technique, called sketching has also been applied in the domain of program synthesis (Solar-Lezama, 2009). Unlike our work, where we use an LLM to suggest templates based on context, the templates in symbolic systems are typically pre-defined or provided in interaction with the user.
Neural Conjecturing and Reinforcement Learning
Recent work on neural conjecturing and reinforcement learning has demonstrated success in specific domains —famously in Euclidian plane geometry through the AlphaGeometry system (Trinh et al., 2024), as well as in reasoning about programs that generate integer sequences from the Online Encyclopedia of Integer Sequences (OEIS) (Gauthier & Urban, 2023, 2025). Poesia et al. (2024) treated conjecturing as a reinforcement learning game in simple propositional logic, arithmetic and group theory, with well formed conjectures generated neurally via constrained decoding. In contrast, we target the broad range of theories represented in proof assistant libraries, and want to avoid domain specific learning by generating generic templates as an intermediate step. The STP system (Dong & Ma, 2025), performs a type of conjecture generation by producing variants of a given seed-statement, with the aim of generating additional training data for a self-playing theorem prover. Lemmanaid, on the other hand, does not rely on having an available seed-statement to start from, and also targets a different use-case, namely to make suggestions to a human doing a formalization.
@@ 3 The Lemmanaid Approach
We have implemented a tool, Lemmanaid, for template-based conjecturing in Isabelle/HOL. The motivation comes from the observation that many lemmas in formalizations share a similar high-level structure and are analogous to each other (Buchberger et al., 2006; Einarsdóttir et al., 2022). The goal of Lemmanaid is to aid a proof assistant in identifying such analogies. We envision a user working on a new mathematical formalization having defined some functions, types and other concepts, and perhaps a few theorems about those. This collection of definitions, which we refer to as a (partial) theory, serves as input to the conjecturing system. Lemmanaid then outputs conjectures that are likely to be useful in this context, allowing the user to make progress in their formalization, or better understand the behavior of the theory they have defined so far (see Figure 1). This happens in two stages: First (neural part), the partial theory is given as input to an LLM which outputs templates likely to be relevant. Secondly (symbolic part), Lemmanaid searches over possible instantiation of those templates in the current theory, to produce concrete conjectures.
Figure 1: High-level overview of Lemmanaid.
Section 3.1 introduces the notion of template-based conjecturing, our template language, and how templates capture information about families of analogous lemmas. We also describe the symbolic engine of Lemmanaid and how it instantiates templates to produce conjectures. Section 3.2 discusses how we train Lemmanaid’s neural engine, a fine-tuned LLM that predicts templates, which are subsequently fed to the symbolic engine. Section 3.3 details direct lemma conjecturing baselines that we implement for Isabelle. Finally, Section 3.4 describes the datasets we use for training and evaluation.
3.1 Template-based conjecturing via Lemmanaid’s Symbolic Engine
@@
?$H_{1}\ x_{1}\ (?H_{2}\ x_{2}\ x_{3})=\ ?H_{2}\ (?H_{1}\ x_{1}\ x_{2})\ (?H_{1}\ x_{1}\ x_{3})$ (2)
?$H_{1}\ x_{1}\ (?H_{1}\ x_{2}\ x_{3})=\ ?H_{1}\ (?H_{1}\ x_{1}\ x_{2})\ x_{3}$ (3)
Note how the function symbols * and $+$ operating on octonions have been replaced by the holes ?$H_{1}$ and ?$H_{2}$, and the variables $a,b,c,x,y$ standardized to $y_{0},y_{1},x_{1},x_{2},x_{3}$. Note that logical symbols, such as negation $\neg$, universal quantifier $\forall$ and equals sign $=$ remain in the template structure. Technically, templates are implemented as instances of Isabelle/HOL’s term datatype [^4] See section 2.2 of the Isabelle/Isar Implementation Manual https://isabelle.in.tum.de/dist/Isabelle2025/doc/implementation.pdf. Our template language represents analogous lemma statements, that have been abstracted and normalized in the following way: Function symbols are replaced with a hole, represented as ?$H_{k}$ where $k$ is a positive integer label. Note that the type of the hole is also an abstraction of the type of the original symbol, with concrete types replaced by type variables. In our examples, the holes in the templates above have types that match any function with two arguments. Each occurrence of a particular function symbol is replaced by the same hole label, even in the case of polymorphic functions where different occurrences of the symbol may have different types. Variables are renamed to $x_{k}$ (or $y_{k}$ in the case of bound variables), where $k$ is a non-negative integer label, and their types set to match those of the corresponding holes. As mentioned, a small set of generic logic symbols have been specified to remain in the template. This is to avoid over-generalization and allow templates to act as representatives for meaningful families of analogous lemmas. See Appendix A.1 for details about these symbols.
@@
We can instantiate or fill the holes in a template by any operator with a matching type. For instance, template 1 can be instantiated using any binary operator, such as subtraction on real numbers or concatenation on lists to obtain a new conjecture. If the operator indeed is non-commutative, we obtain a correct lemma analogous to the lemma octo_product_noncommutative. In this way, templates are generalizations of lemma statements: each template matches many different lemmas that may apply in different theories.
Lemmanaid’s symbolic engine takes as input a template and a list of relevant function names from the current theory. It then searches over the possible instantiations of the holes in the template using the given functions (note that unlike holes the variables are fixed at instantiation, and not replaced by any further terms). Each indexed hole is replaced by one of the given function symbols while ensuring that the resulting candidate lemma is well typed. This may lead to a number of different candidate lemmas being produced. While using templates restricts the search, we still also set a time-out to deal with rare cases where the template contains very many holes or we have a large number of potential functions to instantiate them with.
Revisiting our example, consider again template 3 (associativity). Suppose we give this template to Lemmanaid’s symbolic engine, along with the functions $+$ (addition), -- (subtraction), $sin$, $cos$ and $\;\hat{}\;$ (exponentiation) for real numbers, and the functions $len$ (the length of a list), $rev$ (reversing a list) and @@@@ (concatenate two lists). In the hole-filling step, it will come up with the following candidates:
$x_{1}+x_{2}+x_{3}$ $=x_{1}+(x_{2}+x_{3})$
@@
Note how $sin$, $cos$, $len$, and $rev$ cannot be used to fill the hole ?$H_{1}$, since ?$H_{1}$ is applied to two arguments and therefore requires a binary operator. Also note that only the instantiations with $+$ and @@@@ result in valid lemmas, while the other two conjectures are false. The user can for instance apply Isabelle’s counterexample checker to easily identify these.
3.2 Template Prediction via Lemmanaid’s Neural Engine
Lemmanaid’s neural engine uses a fine-tuned LLM to predict templates given a theory context. Template predictions are inputted to Lemmanaid’s symbolic engine (§3.1). For fine-tuning, we create labeled training data (input-output pairs). Following §3.1, we extract the template of each lemma appearing in a corpus of human-written theory files, and use a string representation of the template as the desired output. For each output, its corresponding input is the list of symbols appearing in the body of the original lemma, as well as contextual information about the symbols. For a given symbol, both its type and any known associated definition convey information about how the symbol can be used. Although definitions provide more complete descriptions of symbols, types are more succinct and they often provide sufficient information to form well-typed templates and lemmas. Section 4.2 evaluates the effectiveness of types and definitions as inputs. Revisiting the examples about octonions, the lemma octo_distrib_left would result in the datapoint of the shape:
Input: [Symbols: *, +] [Types of *, +] [Defs of *, +]; Output: [?$H_{1}\ x_{1}\ (?H_{2}\ x_{2}\ x_{3})=\ ?H_{2}\ (?H_{1}\ x_{1}\ x_{2})\ (?H_{1}\ x_{1}\ x_{3})$]

### v5

@@ Abstract
Mathematicians and computer scientists are increasingly using proof assistants to formalize and check correctness of complex proofs. This is a non-trivial task in itself, however, with high demands on human expertise. Can we lower the bar by introducing automation for conjecturing helpful, interesting and novel lemmas? We present the first neuro-symbolic lemma conjecturing tool, Lemmanaid, designed to discover conjectures by drawing analogies between mathematical theories. Lemmanaid uses a fine-tuned LLM to generate lemma templates that describe the shape of a lemma, and symbolic methods to fill in the details. We compare Lemmanaid against the same LLM fine-tuned to generate complete lemma statements (a purely neural method), as well as a fully symbolic conjecturing method. Lemmanaid consistently outperforms both neural and symbolic methods on test sets from Isabelle’s HOL library and from its Archive of Formal Proofs (AFP). Using DeepSeek-coder-6.7B as a backend, Lemmanaid discovers 50% (HOL) and 28% (AFP) of the gold standard reference lemmas, 8-13% more than the corresponding neural-only method. Ensembling two Lemmanaid versions with different prompting strategies further increases performance to 55% and 34% respectively. In a case study on the formalization of Octonions, Lemmanaid discovers 79% of the gold standard lemmas, compared to 62% for neural-only and 23% for the state of the art symbolic tool. Our result show that Lemmanaid is able to conjecture a significant number of interesting lemmas across a wide range of domains covering formalizations over complex concepts in both mathematics and computer science, going far beyond the basic concepts of standard benchmarks such as miniF2F, PutnamBench and ProofNet.
Lemmanaid: Neuro-Symbolic Lemma Conjecturing
Yousef Alhessi1, Emily First & Sorin Lerner Department of Computer Science and Engineering University of California, San Diego, USA {yalhessi,emfirst, lerner}@ucsd.edu
Sólrún Halla Einarsdóttir1 , George Granberry, Moa Johansson & Nicholas Smallbone Department of Computer Science and Engineering Chalmers University of Technology & University of Gothenburg Gothenburg, Sweden {slrn,georgegr,moa.johansson,nicsma}@chalmers.se 1Joint first authorship
@@ 1 Introduction
Learning to construct new, interesting, and useful lemmas is an important and long standing challenge in AI for mathematical reasoning Lenat (1976); Fajtlowicz (1988); Ireland and Bundy (1996); Colton (2002), yet it remains an underexplored area (Yang et al., 2025). Automatically discovered lemmas can both aid a human user working on a mathematical formalization and strengthen automated theorem provers. In this work, we examine how the generalization capabilities of LLMs can be combined with symbolic tools, to generate lemmas by a process of analogy (Figure 1). The LLM suggests generic templates describing learned families of analogous lemmas, which are then instantiated to concrete theories symbolically. Previous work has shown that such families of analogous lemmas indeed occur in proof assistant libraries Einarsdóttir et al. (2022); Heras et al. (2013).
Figure 1: High-level overview of Lemmanaid.
Proof assistants like Lean de Moura et al. (2015) and Isabelle Nipkow et al. (2002) are becoming increasingly important among research mathematicians who want complex proofs checked for guaranteed correctness. Following the formalization of Kepler’s conjecture Hales et al. (2017), there have been several projects formalizing recent results in research mathematics Scholtze (2020); Commelin and Topaz (2024); Gowers et al. (2023); Tao (2023). However, formalizing mathematics is a non-trivial task, which in turn has sparked research in areas where large language models might be of assistance: primarily autoformalization, translating informal proofs to formal language Wang et al. (2018); Szegedy (2020); Wu et al. (2022), and formal proof synthesis, suggesting the next steps for the proof assistant Polu and Sutskever (2020); Jiang et al. (2022, 2023); First et al. (2023). Here, the LLM and proof assistant complement each other: LLMs can provide a more flexible and powerful proof search while hallucinations in proofs are caught by the proof assistant and easily discarded. We address a third challenge, conjecturing, with the aim of providing a first tool, Lemmanaid, towards generic conjecturing over a broad range of mathematical theories in domains from realistic formalizations by researchers in computer science and mathematics. We move beyond the basic domains appearing in standard benchmarks for LLM-based mathematics reasoning, which have been derived from high school and undergraduate level mathematics competitions and standard textbooks, like miniF2F Zheng et al. (2022), PutnamBench Tsoukalas et al. (2024) and ProofNet Azerbayev et al. (2023).
Conjecturing: Methods and Evaluation
@@
A weakness of neural conjecturing is that it sometimes generates repetitive, redundant or trivial lemmas, or hallucinates undefined symbols in the formalization (Urban and Jakubův, 2020; Rabe et al., 2021; Johansson and Smallbone, 2023). Scaling the model is a commonly attempted remedy, albeit computationally expensive and most effective in domains where background libraries have already been formalized, such as the theories needed for e.g. math olympiad problems Zheng et al. (2022). Symbolic methods, on the other hand, can be designed and programmed to avoid repetition and redundancy. Previous symbolic tools  (Smallbone et al., 2017; Einarsdóttir et al., 2021; Singher and Itzhaky, 2021) have been used to successfully discover, for example, lemmas needed in automated (co-)inductive provers (Johansson et al., 2014; Einarsdóttir et al., 2018, 2024; Kurashige et al., 2024) over (co-)recursive datatypes. However, symbolic tools are limited in the shape, size and domain of lemmas they can generate, and do not scale well to larger sets of inputs. To address these shortcomings, we propose a novel neuro-symbolic lemma conjecturing approach and tool: Lemmanaid. In Lemmanaid, an LLM is trained to generate lemma templates that describe the shape of a family of analogous lemmas, rather than directly generating complete lemmas (see Figure 1). Symbolic synthesis methods are then used to fill in the details. In this way, we leverage the best of both neural and symbolic methods. The LLM suggests appropriate analogous lemma-patterns likely to be relevant for the theory at hand. The symbolic engine ensures correctness and novelty, while keeping the search space manageable. As far as we are aware, this is the first work focusing on neuro-symbolic lemma conjecturing.
Conjecturing can broadly be divided into two main methodologies, both of which both have been attempted with both symbolic and neural methods Johansson (2019): Top-down methods find relevant lemmas for a specific proof attempt Ireland and Bundy (1996); Heras et al. (2013). This includes recent work using test time reinforcement learning, where custom curricula are produced to train a specialized neural prover Dong and Ma (2025); Hubert et al. (2025); Chen et al. (2025). Bottom-up methods, to which our work belongs, instead aim to find interesting lemmas from a set of definitions, without any pre-defined top-level goal Lenat (1976); McCasland et al. (2017); Urban and Jakubův (2020); Johansson and Smallbone (2023). Evaluation metrics are typically different for top-down and bottom-up approaches. For top-down conjecturing, increased proof success rate (of original proofs) is a common measure, while for bottom-up methods, like ours, a notion of interestingness of conjectures must be defined. Using human-written formalizations as a gold-standard is a common evaluation strategy Montano-Rivas et al. (2012); Smallbone et al. (2017); Einarsdóttir et al. (2018); Urban and Jakubův (2020). For Lemmanaid, we therefore measure the coverage of (unseen test-set) lemmas that can be discovered by Lemmanaid for (a subset of) Isabelle’s HOL library [^1] https://isabelle.in.tum.de/dist/library/HOL/index.html and from its Archive of Formal Proofs (AFP) [^2] https://www.isa-afp.org.
Contributions
In our experiments, we use DeepSeek-coder-1.3b and 6.7b, and Llama-3.2-1b language models. For computational resource reasons we could not evaluate larger models. We show that the Lemmanaid approach compares favorably to the results achievable using purely neural or purely symbolic conjecturing. The main contributions of our work are:
@@
  • Lemmanaid, the first neuro-symbolic lemma conjecturing approach that uses an LLM to suggest templates and a symbolic engine to instantiate templates as candidate lemmas.
@@
  • An broad evaluation of Lemmanaid on the Isabelle proof assistant’s HOL and AFP libraries. This goes beyond evaluations of prior tools that have focused on specific mathematical domains.
@@ 2 Related Work
Proof Assistants, such as Isabelle (Nipkow et al., 2002) and Lean (de Moura et al., 2015), are increasingly being used to check proofs in both mathematics and computer science for correctness. To do so, the user needs to formalize their theory, by translating it into the formal language of the proof assistant. They then interact with the system to construct a proof by stringing together calls to tactics, each executing and checking some part of the proof. Popular tools like Sledgehammer can automate (parts of) many proofs by selecting a suitable set of previously proved lemmas, and sending the conjecture to an automated first-order prover or SMT-solver (Blanchette et al., 2011).
Formalizing theories in proof assistants is a non-trivial task, which has sparked interest in autoformalization: translating definitions, theorems and lemmas written in natural language to the formal language of a proof assistant (Wang et al., 2018; Szegedy, 2020; Wu et al., 2022). Recent work by Zhang et al. (2025) shows that autoformalization on realistic mathematical definitions and problems, appearing on ArXiV and Wikipedia, is still a challenge for LLMs. They present a new benchmark for autoformalization of mathematical definitions, going beyond the basic definitions of e.g. miniF2F. We see Lemmanaid as an excellent complement to this kind of work: autoformalized definitions could be passed on to Lemmanaid to provide a richer initial formalization.
Proof Synthesis
@@
Even with definitions and statements formalized, constructing the required proofs from various tactics is again non-trivial, even with the help of tools like Sledgehammer. This has motivated work on various LLM-driven methods for synthesizing proof scripts for proof assistants such as Isabelle or Lean, either step by step, whole proofs at once, or via proof sketches (Polu and Sutskever, 2020; Wang et al., 2024; Ren et al., 2025; Hubert et al., 2025; Chen et al., 2025; Wang et al., 2025; Lin et al., 2025; Jiang et al., 2022; First et al., 2023; Jiang et al., 2023). Our work does not focus on producing proofs, but on suggesting suitable conjectures similar to the kind of lemmas that humans write down in libraries of formal proofs, which could in the future be combined with, and complement, proof synthesis systems.
Neural Conjecturing and Reinforcement Learning
Recent top-down conjecturing methods have successfully been used in test time reinforcement learning as a means to generate additional data with the aim of creating a curriculum for training a neural prover Dong and Ma (2025); Hubert et al. (2025); Chen et al. (2025). This has contributed to the recent gold medal performances on International Math Olympiad problems Hubert et al. (2025); Chen et al. (2025). Here, conjectures are typically variants of some given seed statement of interest. Lemmanaid, on the other hand, uses a bottom-up conjecturing approach and targets a different use-case. It aims to make novel suggestions directly from definitions, interesting to a human doing a formalization, rather than generating variants of a given statement.
Other methods combining neural conjecturing and reinforcement learning have demonstrated success in specific domains, such as inductive proofs about synthesized programs that generate integer sequences (Gauthier and Urban, 2025). The LEGO prover attempts to introduce intermediate proof statements as reusable lemmas Wang et al. (2024) for other problems in the same domain. Poesia et al. (2024) treat conjecturing as a reinforcement learning game in simple propositional logic, arithmetic and group theory, with well formed conjectures generated neurally via constrained decoding. Tsoukalas et al. (2025) introduce a system combining symbolic rules and reinforcement learning to form conjectures in elementary number theory and on finite fields. The system is evaluated against a gold-standard set of conjectures, while simultaneously developing an interestingness function over conjectures. In contrast, Lemmanaid targets the broad range of theories from research in computer science and mathematics represented in proof assistant libraries, and aims to avoid domain specific learning by generating generic templates as an intermediate step.
@@
Following the observation that many mathematical theories share analogical lemmas of similar shapes, Buchberger et al. (2006) first proposed to use templates as human-provided guidance in mathematical theory exploration, by conjecturing interesting lemmas by analogy to known shapes. This has been implemented in a range of symbolic lemma conjecturing systems, including some targeting Isabelle/HOL (Montano-Rivas et al., 2012; Heras et al., 2013; McCasland et al., 2017; Einarsdóttir et al., 2021; Nagashima et al., 2023). A similar technique called sketching has also been applied in the domain of program synthesis (Solar-Lezama, 2009). In the above works, templates have typically been provided by the human user. Nye et al. (2019) used a neural network to propose program sketches which were then filled in by a symbolic program synthesizer, thus sharing some features with our work.
@@ 3 The Lemmanaid Approach
We have implemented a tool, Lemmanaid, for template-based conjecturing in Isabelle/HOL. The motivation comes from the observation that many lemmas in formalizations share a similar high-level structure and are analogous to each other (Buchberger et al., 2006; Einarsdóttir et al., 2022). The goal of Lemmanaid is to aid a proof assistant in identifying such analogies. We envision a user working on a new mathematical formalization; having defined some functions, types and other concepts, and perhaps a few theorems about these. This collection of definitions, which we refer to as a (partial) theory, serves as input to the conjecturing system. Lemmanaid then outputs conjectures that are likely to be useful in this context, allowing the user to make progress in their formalization, or better understand the behavior of the theory they have defined so far (see Figure 1). This happens in two stages: First (neural part), the partial theory is given as input to an LLM which outputs templates likely to be relevant. Second (symbolic part), Lemmanaid searches over possible instantiations of those templates in the current theory, to produce concrete conjectures.
We make all code, experimental scripts, data, and models publicly available online [^3] https://anonymous.4open.science/r/Lemmanaid-ACL. See Appendix B for details about computing resources and replication parameters used in our evaluation.
3.1 Lemmanaid’s Symbolic Engine
@@
$H_{1}\ x_{1}\ (?H_{2}\ x_{2}\ x_{3})=\\ ?H_{2}\ (?H_{1}\ x_{1}\ x_{2})\ (?H_{1}\ x_{1}\ x_{3})$ (2)
$H_{1}\ x_{1}\ (?H_{1}\ x_{2}\ x_{3})=\ ?H_{1}\ (?H_{1}\ x_{1}\ x_{2})\ x_{3}$ (3)
Note how the function symbols * and ++ operating on octonions have been replaced by the holes $H_{1}$ and $H_{2}$, and the variables $a,b,c,x,y$ standardized to $y_{0},y_{1},x_{1},x_{2},x_{3}$. Note that logical symbols, such as negation $\neg$, universal quantifier $\forall$ and equals sign == remain in the template structure. Technically, templates are implemented as instances of Isabelle/HOL’s term datatype [^4] See section 2.2 of the Isabelle/Isar Implementation Manual https://isabelle.in.tum.de/dist/Isabelle2025/doc/implementation.pdf. Our template language represents analogous lemma statements, that have been abstracted and normalized in the following way: Function symbols are replaced with a hole, represented as $H_{k}$ where $k$ is a positive integer label. Note that the type of the hole is also an abstraction of the type of the original symbol, with concrete types replaced by type variables. In our examples, the holes in the templates above have types that match any function with two arguments. Each occurrence of a particular function symbol is replaced by the same hole label, even in the case of polymorphic functions where different occurrences of the symbol may have different types. Variables are renamed to $x_{k}$ (or $y_{k}$ in the case of bound variables), where $k$ is a non-negative integer label, and their types set to match those of the corresponding holes. As mentioned, a small set of generic logic symbols have been specified to remain in the template. This is to avoid over-generalization and allow templates to act as representatives for meaningful families of analogous lemmas. See Appendix A for details about these symbols.
@@
We can instantiate or fill the holes in a template by any operator with a matching type. For instance, template 1 can be instantiated using any binary operator, such as subtraction on real numbers or concatenation on lists, to obtain a new conjecture. If the operator indeed is non-commutative, we obtain a correct lemma analogous to the lemma octo_product_noncommutative. In this way, templates are generalizations of lemma statements: each template matches many different lemmas that may apply in different theories.
Lemmanaid’s symbolic engine takes as input a template and a list of relevant function names from the current theory. It then searches over the possible instantiations of the holes in the template using the given functions (note that unlike holes the variables are fixed at instantiation, and not replaced by any further terms). Each indexed hole is replaced by one of the given function symbols while ensuring that the resulting candidate lemma is well typed. This may lead to a number of different candidate lemmas being produced. While using templates restricts the search, we still also set a time-out to deal with rare cases where the template contains very many holes or we have a large number of potential functions to instantiate them with. We also experimented with using an LLM to also instantiate the templates, but found it performed worse than the symbolic engine as well as incurring a greater computational cost (see Appendix C).
Revisiting our example, consider again template 3 (associativity). Suppose we give this template to Lemmanaid’s symbolic engine, along with the functions ++ (addition), - (subtraction), sinsin, coscos and $\hat{}\;$ (exponentiation) for real numbers, and the functions lenlen (the length of a list), revrev (reversing a list) and @@ (concatenate two lists). In the hole-filling step, it will come up with the following candidates:
$(x_{1}+x_{2})+x_{3}$ $=x_{1}+(x_{2}+x_{3})$
@@
Note how sinsin, coscos, lenlen, and revrev cannot be used to fill the hole $H_{1}$, since $H_{1}$ is applied to two arguments and therefore requires a binary operator. Also note that only the instantiations with ++ and @@ result in valid lemmas, while the other two conjectures are false. The user can for instance apply Isabelle’s counterexample checker to easily identify these.
3.2 Lemmanaid’s Neural Engine
Lemmanaid’s neural engine uses a fine-tuned LLM to suggest templates given a theory context, which are then passed to Lemmanaid’s symbolic engine (§3.1). For fine-tuning, we create labeled training data (input-output pairs) following §3.1: we extract the template of each lemma appearing in a corpus of human-written theory files, and use a string representation of the template as the desired output. For each output template, its corresponding input is the list of symbols appearing in the body of the original lemma, as well as contextual information about the symbols. For a given symbol, both its type and any known associated definition convey information about how the symbol can be used. Although definitions provide more complete descriptions of symbols, types are more succinct and they often provide sufficient information to form well-typed templates and lemmas. Section 4.2 evaluates the effectiveness of types and definitions as inputs. Revisiting the examples about octonions, the lemma octo_distrib_left would result in the data-point of the shape:
Input: [Symbols: *, +] [Types of *, +] [Defs of *, +];
Output: $H_{1}\ x_{1}\ (?H_{2}\ x_{2}\ x_{3})=\\ ?H_{2}\ (?H_{1}\ x_{1}\ x_{2})\ (?H_{1}\ x_{1}\ x_{3})$]